$SPAD/src/input graphics.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input for page OneVariableGraphicsExamplePage
)clear all
draw(sin tan x - tan sin x, x = 0..6)
draw(sin x + cos x, x = 0..2*%pi)
draw(sin(1/x), x = -1..1)
draw(x * sin(1/x), x = -1..1)
All
-- Input for page ViewportPage
)clear all
v := draw(x*x-y*y,x=-1..1,y=-1..1)
write(v,"saddle","pixmap")
-- Input for page ImplicitCurveGraphicsExamplePage
)clear all
draw(x * y = 1, x, y, range == [-3..3, -3..3])
draw(y**2 + y = x**3 - x, x, y, range == [-2..2, -2..1])
p := ((x**2 + y**2 + 1) - 8*x)**2 - (8*(x**2 + y**2 + 1) - 4*x - 1)
draw(p = 0, x, y, range == [-1..11, -7..7], title == "Cartesian Ovals")
q := (x**2 + y**2 + 7**2)**2 - (6**4 + 4*7**2*x**2)
draw(q = 0, x, y, range == [-10..10, -4..4], title == _
  "Cassinian oval with two loops")
All
-- Input for page OneVariableGraphicsPage
)clear all
draw(sin(tan(x)) - tan(sin(x)),x = 0..6)
draw(sin(tan(x)) - tan(sin(x)),x = 10..16,title == "y = sin tan x - tan sin x")
f(x) == (x-1)*(x-2)*(x-3)
draw(f, 0..2, title == "y = f(x) on [0,2]")
draw(f, 0..4,title == "y = f(x) on [0,4]")
-- Input for page PolarGraphicsExamplePage
)clear all
draw(1,t = 0..2*%pi, coordinates == polar)
draw(t,t = 0..100, coordinates == polar)
draw(sin(4*t), t = 0..2*%pi, coordinates == polar)
draw(2 + 3 * sin t, t = 0..2*%pi, coordinates == polar)
All
-- Input for page ParametricCurveGraphicsExamplePage

)clear all
draw(curve(cos(t/(1+sin(t)**2)), sin(t)*cos(t)/(1+sin(t)**2)), t = -%pi..%pi)
draw(curve(9*sin(3*t/4), 8*sin(t)), t = -4*%pi..4*%pi)
draw(curve(sin(t)*sin(2*t)*sin(3*t), sin(4*t)*sin(5*t)*sin(6*t)),t = 0..2*%pi)
draw(curve(cos(4*t)*cos(7*t), cos(4*t)*sin(7*t)), t = 0..2*%pi)
All
-- Input for page PolarGraphicsPage
)clear all
draw(1,t = 0..2*%pi,coordinates == polar, title == "The Unit Circle")
draw(sin(17*t), t = 0..2*%pi, coordinates == polar, title == "A Petal Curve")
f(t) == cos(4*t/7)
draw(f, t = 0..2*%pi, coordinates == polar)
draw(f, t = 0..14*%pi, coordinates == polar)
-- Input for page ParametricCurveGraphicsPage
)clear all
draw(curve(sin(t)*sin(2*t)*sin(3*t), sin(4*t)*sin(5*t)*sin(6*t)), t = 0..2*%pi)
draw(curve(cos(t), sin(t)), t = 0..2*%pi, title == "The Unit Circle")
f(t) == sin(3*t/4)
g(t) == sin(t)
draw(curve(f,g), 0..%pi)
draw(cruve(f,g) ,%pi..2*%pi)
draw(curve(f,g), -4*%pi..4*%pi)
-- Input for page AssortedGraphicsExamplePage
)clear all
draw(sin(x * y), x = -2.5..2.5, y = -2.5..2.5)
draw(sin tan x - tan sin x,x = 0..6)
draw(curve(sin(t)*sin(2*t), sin(3*t)*sin(4*t)), t = 0..2*%pi)
draw(curve(sin(t)*sin(2*t), sin(3*t)*sin(4*t), sin(5*t)*sin(6*t)), t = 0..2*%pi)
draw(sin(17*t), t = 0..2*%pi, coordinates == polar)
draw(y**2 + y = x**3 - x, x, y,range == [-2..2,-2..1])
All

———-

References^{[1] nothing}

5